# Evo-0: Vision-Language-Action Model with Implicit Spatial Understanding

### v1

@@ 4.2 Task Setup
We design five tasks for real-world robot evaluation, which span a range of spatial understanding challenges, from fine-grained geometric alignment to pick-and-place and transparent object interaction. In particular, each task has a low tolerance for spatial error, as minor inaccuracies in the spatial predictions can lead to task failure. This makes them well-suited for assessing if representations from VGGT can enhance VLA’s spatial understanding.
Figure 2: Illustration of the task setup.
@@
    Middle bottle grasping. Three bottles are closely placed in a row, and the robot is instructed to pick the middle one. This setup mimics a grocery store scenario, where items are densely arranged on shelves. Success is defined as picking up the middle bottle without touching or knocking over the adjacent ones.
@@
Figure 3: Qualitative results of our model in real-world tasks.
4.3 Real-World Experiment
@@
Table 1: Success rates for five real-world tasks. The number in parentheses represents the number of trails for each task. In particular, Task 1 is evaluated using scores; for example, 68.67% denotes that our model achieves an average score of 3.43 out of a maximum of 5.
Quantitative Results. We evaluate the effectiveness of our proposed method by comparing it against the baseline model $\pi_{0}$. The quantitative results are presented in the Table 1. Across all tasks, our method achieves consistent improvements over the baseline, indicating that the implicit 3D geometry features contribute positively to task performance. Notably, our model demonstrates the largest performance gain on Task 2 (peg-in-hole insertion), a particularly challenging task that demands accurate spatial reasoning. Furthermore, Task 3 (middle bottle grasping) poses a substantial challenge due to the narrow margin between adjacent bottles, requiring the gripper to perform careful, collision-free insertion and grasping. Our method exhibits reasonable improvement on this task compared to the baseline, demonstrating enhanced spatial understanding and control in cluttered environments. Overall, we achieve a 28.88% performance gain in the average success rate.
@@
5 Conclusion
@@
In this paper, we explore using implicit 3D representations to enhance spatial understanding in Vision-Language-Action (VLA) models. By leveraging features from the Visual Geometry Grounded Transformer (VGGT), trained on large-scale 2D–3D paired data, we inject strong geometric priors into VLA models without relying on explicit 3D inputs. Through extensive experiments across five spatially challenging tasks, we demonstrate that our approach significantly outperforms baseline models, validating the effectiveness of the proposed implicit geometric prior integration. Our method offers a simple and efficient solution for enhancing spatial understanding in VLA systems.

### v2

@@ IV-C Real-World Experiments
Task Setup. We design five tasks for real-world robot evaluation, which span a range of spatial understanding challenges, from fine-grained geometric alignment to pick-and-place and transparent object interaction. In particular, each task has a low tolerance for spatial error, as minor inaccuracies in the spatial predictions can lead to task failure. This makes them well-suited for assessing if representations from VGGT can enhance VLA’s spatial understanding.
@@
    Middle bottle grasping. Three bottles are closely placed in a row, and the robot is instructed to pick the middle one. This setup mimics a grocery store scenario, where items are densely arranged on shelves. Success is defined as picking up the middle bottle without knocking over the adjacent ones.
@@
Implementation Details. For each task, we collect 100 expert demonstrations using tele-operation. To promote diversity and robustness, the positions of objects and targets are randomly perturbed during data collection. In our real-world experiments, we keep the training and parameter settings identical to those used in simulation experiments.
@@
TABLE I: Success rates for five real-world tasks. The number in parentheses represents the number of trials for each task. In particular, Task 1 is evaluated using scores; for example, 68.67% denotes that our model achieves an average score of 3.43 out of a maximum of 5.
Quantitative Results. We evaluate the effectiveness of our proposed method by comparing it against the baseline model $\pi_{0}$. The quantitative results are presented in the Table I. Across all tasks, our method achieves consistent improvements over the baseline, indicating that the implicit 3D geometry features contribute positively to task performance. Notably, our model demonstrates the largest performance gain on Task 2 (peg-in-hole insertion), a particularly challenging task that demands accurate spatial reasoning. Furthermore, Task 3 (middle bottle grasping) poses a substantial challenge due to the narrow margin between adjacent bottles, requiring the gripper to perform careful, collision-free insertion and grasping. Our method exhibits reasonable improvement on this task compared to the baseline, demonstrating enhanced spatial understanding and control in cluttered environments. Overall, we achieve a 28.88% performance gain in the average success rate.
Figure 5: Qualitative results in real-world tasks. Detailed execution sequences of the five real-world evaluation tasks are visualized, illustrating the step-by-step progression for each scenario.
@@
IV-D Robustness Experiments
Task Setup. In the robustness experiments, we begin with a simplified version of real-world Task 4 and progressively introduce additional disturbances to assess the model’s robustness. In this task, the robot is required to pick up a beverage can and place it onto a designated location on a shelf. Compared with the original Task 4 configuration, we apply two modifications to construct a base scenario: (i) the camera viewpoint is changed from a side view to a frontal view, mitigating visual occlusions during manipulation; and (ii) the shelf position is adjusted to facilitate easier placement, thereby reducing the likelihood of unintended collisions. Under this simplified condition, both $\pi_{0}$ and Evo-0 attain identical success rates. However, as disturbances are gradually introduced, Evo-0 consistently exhibits higher robustness than $\pi_{0}$.
Disturbance Conditions. To systematically evaluate robustness, we design five categories of disturbance conditions, which are illustrated in Figure 6: (i) introduction of an unseen distractor object, (ii) variation in background color, (iii) displacement of the target position, (iv) change in target height, and (v) shift in camera viewpoint. All conditions lie outside the training distribution. Each disturbance targets a distinct aspect of generalization, enabling a comprehensive assessment of the model’s robustness to environmental variation and unseen distractors.
Figure 6: Distractor settings of robustness experiments. We visualize five disturbance conditions used in the robustness experiments: unseen distractor object, background color variation, target position variation, target height variation, and camera viewpoint variation. In the top right corner, we also present the execution sequence of a successful trial.
Quantitative Results. We conduct comprehensive evaluations across all five disturbance settings, with the quantitative results summarized in Table II.
  1. 1.
    Unseen distractor object. In the unseen distractor object experiment, the target can and a distractor bottle are placed on opposite sides of the shelf. The task requires the robot to correctly identify the can, grasp it, and place it at the designated location. This setting evaluates both the model’s ability to distinguish the correct object and to execute precise placement in the presence of visual distractors. Experimental results show that Evo-0 substantially outperforms $\pi_{0}$ on both aspects. In the object selection phase, Evo-0 achieves a 100% success rate, representing a 40 percentage point improvement over $\pi_{0}$ (60%). For the full grasp-and-place sequence, Evo-0 attains a 70% success rate, outperforming $\pi_{0}$ (20%) by 50 percentage points. These results demonstrate that Evo-0 exhibits stronger robustness in both object discrimination and manipulation accuracy.
  2. 2.
    Background color variation. In the background color variation experiment, the tablecloth is changed to modify the scene appearance. This setting evaluates how each model maintains placement accuracy under background changes. Experimental results show that Evo-0 achieves approximately 10% higher accuracy than $\pi_{0}$ in this condition. These findings indicate that Evo-0 possesses stronger robustness to variations in background appearance.
  3. 3.
    Target position variation. In the target position variation experiment, the mat is shifted 10 mm, 20 mm, and 30 mm backward along the shelf across separate test groups. This setting assesses robustness to spatial displacement of the target. Evo-0 consistently achieves approximately 10% higher success rates than $\pi_{0}$ under the 20 mm and 30 mm shifts, indicating stronger robustness and generalization to positional perturbations.
  4. 4.
    Target height variation. To assess robustness to vertical displacement, the thickness of the target mat is increased by 5 mm, 10 mm, and 15 mm across separate test groups. As height increases, both models exhibit performance degradation. However, Evo-0 consistently outperforms $\pi_{0}$, with accuracy margins of 20% and 10% under the 10 mm and 15 mm conditions, respectively. These results indicate that Evo-0 demonstrates stronger robustness and enhanced spatial reasoning under vertical perturbations.
  5. 5.
    Camera viewpoint variation. In this setting, the wrist camera remains fixed, while the base camera is tilted upward and downward by 10° in separate experiments to introduce novel camera viewpoints. This variation evaluates robustness to changes in camera perspective. Both models experience performance degradation under these shifts; however, Evo-0 consistently outperforms $\pi_{0}$, achieving 20% and 10% higher accuracy in the upward and downward conditions, respectively. These results suggest that Evo-0 demonstrates stronger robustness to camera viewpoint changes.
Condition $\pi_{0}$ Ours
Base 70% 70%
Unseen distractor object
Pick correct rate 60% 100%
Overall correct rate 20% 70%
Background color variation
Add blue tablecloth 60% 70%
Target position variation
10 mm backward 60% 60%
20 mm backward 20% 30%
30 mm backward 10% 20%
Target height variation
5 mm higher 70% 70%
10 mm higher 10% 30%
15 mm higher 0% 10%
Camera viewpoint variation
10 degrees up 40% 60%
10 degrees down 30% 40%
TABLE II: Success rates for robustness experiments. Comparison of success rates between $\pi_{0}$ and Ours under different disturbance conditions in real-world task robustness experiments.
V Conclusion
@@
In this paper, we explore using implicit 3D representations to enhance spatial understanding in Vision-Language-Action (VLA) models. By leveraging features from the Visual Geometry Grounded Transformer (VGGT), trained on large-scale 2D–3D paired data, we inject strong geometric priors into VLA models without relying on explicit 3D inputs. Through extensive experiments across spatially challenging tasks in both simulation and real-world, we demonstrate that our approach significantly outperforms baseline models, validating the effectiveness of the proposed implicit geometric prior integration. Our method offers a simple and efficient solution for enhancing spatial understanding in VLA systems.